OpenStudy (anonymous):

help with Algebra 2?

Nnesha (nnesha):

what you got ??

OpenStudy (zale101):

what happens when you plug in 0 to x?

OpenStudy (zale101):

@nello

OpenStudy (anonymous):

the x value is 0?

OpenStudy (anonymous):

sorry I mean y=3

OpenStudy (zale101):

3+6(0)=? what is y=?

OpenStudy (zale101):

yes

OpenStudy (zale101):

in your table, it says it equals to 2, is that reasonable?

Nnesha (nnesha):

ok so look your table where x is 0 there should be y-intercept

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

oh no no its not

Nnesha (nnesha):

and for the graph where line touch the y-axis there is your y-intercept

OpenStudy (zale101):

y=3+6x is not the answer, the y-intercept is not at (0,3)

OpenStudy (anonymous):

yeah so the answer would be A

OpenStudy (zale101):

y is 2 more than six times x, means y=2+6x ? correct if you plug in 0 to x you'd get (0,2) for the y intercept so A is not the answer

OpenStudy (anonymous):

oh yes okay sorry

OpenStudy (zale101):

we said A was wrong, remember

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

when I plug in the numbers in the table into the equation its wrong

OpenStudy (anonymous):

So I'm guessing the one that doe snot describe the same situation is the table??

Nnesha (nnesha):

no table is right when you use slope formula you should get same slope which is 6x and same y-intercept which is 2 so A is incorrect statement so answer is A

OpenStudy (anonymous):

thanks :)